# MedGR2: Breaking the Data Barrier for Medical Reasoning via Generative Reward Learning

### v1

@@ Abstract
The application of Vision-Language Models (VLMs) in medicine is critically hampered by the scarcity of high-quality, expert-annotated data. Supervised Fine-Tuning (SFT) on existing datasets often leads to poor generalization on unseen modalities and tasks, while Reinforcement Learning (RL), a promising alternative, is stymied by the lack of reliable reward signals in this data-scarce domain. To break this impasse, we introduce Generative Reward Learning for Medical Reasoning (MedGR2), a novel framework that creates a self-improving virtuous cycle. MedGR2 co-develops a data generator and a reward model, enabling the automated, continuous creation of high-quality, multi-modal medical data that serves as both a superior training source for SFT and RL. Our experiments demonstrate that SFT with MedGR2-produced data already surpasses baselines trained on large-scale, human-curated datasets. Crucially, when leveraging this data for RL via Group Relative Policy Optimization (GRPO), our model achieves state-of-the-art cross-modality and cross-task generalization, significantly outperforming specialized RL-based methods. Furthermore, our compact model, empowered by MedGR2, achieves performance competitive with foundation models possessing over 10 times more parameters. MedGR2 presents a new paradigm for data-efficient learning in high-stakes domains, transforming the problem from data scarcity to data generation and unlocking the full potential of RL for building truly generalizable medical AI.
1 Introduction
@@
Vision-language models (VLMs) hold great promise for transforming clinical reasoning, from multimodal diagnosis to radiology report generation. Their ability to jointly process and reason over medical images and texts offers an unprecedented opportunity to improve decision support, reduce clinician burden, and enable generalizable AI in high-stakes domains. Yet this vision is fundamentally bottlenecked by a key obstacle: the scarcity of high-quality, expert-annotated medical data. Compared to general-domain tasks, acquiring medical data is constrained by privacy, annotation cost, and the domain-specific nature of expertise, which results in narrow task coverage and limited generalization across unseen modalities and tasks.
@@
To address this challenge, the research community has explored two dominant paths, each with clear limitations. The first approach focuses on large-scale data curation and supervised fine-tuning (SFT). Some systems(Chen et al. 2024) attempt to clean and reformat noisy biomedical corpora such as PubMed using large vision-language models (e.g., GPT-4O), producing instruction-style datasets for downstream training. While effective at increasing data scale, these methods remain fundamentally limited by the quality of their source material and the domain alignment of the models used for cleaning. More importantly, as recent studies highlight, SFT-trained models often memorize training distributions and struggle with generalization to new modalities and tasks. This constitutes a critical gap for real-world clinical deployment.(Zhao et al. 2024).
@@
The second path is reinforcement learning (RL), which offers a more powerful paradigm for learning generalizable reasoning strategies. For instance, some works(Lai et al. 2025; Pan et al. 2025) demonstrate that training with RL using small, expert-annotated preference datasets significantly improves cross-task and cross-modal performance over SFT baselines. However, RL in medicine faces a structural limitation: it relies on reward signals that are expensive to obtain, requiring human expert judgments for reward modeling or preference annotation. This creates a paradox: RL requires high-quality supervision to improve generalization, but such supervision is itself too costly to scale.
@@ 2 Related Work
Constructing high-quality, large-scale datasets remains a central bottleneck in developing medical VLMs. Manual annotation is prohibitively expensive and labor-intensive, prompting a surge of semi-automated solutions. HuatuoGPT-Vision (Chen et al. 2024) reformats PubMed-derived image-text pairs into instruction-following VQA samples using GPT-4V. LLaVA-Med (Li et al. 2023a) synthesizes multimodal QA pairs by prompting ChatGPT with biomedical figures. PMC-VQA (Zhang et al. 2023) leverages large language models and rule-based heuristics to mine high-quality triplets from scientific publications. Flamingo-CXR (Tanno et al. 2025) demonstrates the potential of radiologist-validated datasets, yet the extensive human oversight required severely limits scalability. While these approaches offer valuable contributions, they remain constrained by source data noise, domain coverage gaps, and alignment inconsistencies. In contrast, our framework abandons static corpora altogether, employing a self-improving generative pipeline to synthesize clinically meaningful, task-aligned multimodal instruction data from scratch.
@@
To mitigate data scarcity, generative approaches such as GANs and diffusion models have been widely adopted for synthesizing realistic medical images (Yi, Walia, and Babyn 2019; Croitoru et al. 2023; Li and Guo 2025). While effective for augmenting training sets in pixel-level tasks such as classification or segmentation, these techniques fall short in producing semantically coherent, instruction-following multimodal pairs required for medical reasoning. State-of-the-art VLMs demand data that captures clinical logic, procedural knowledge, and image-language alignment. Our framework transcends pixel-level synthesis by generating complete reasoning tasks paired with reward-guided quality filtering, enabling scalable construction of instruction-grade datasets for medical VLMs.
@@
Although SFT has facilitated early progress in VLMs, it often encourages rote pattern learning over robust generalization (Chu et al. 2025). This limitation is especially detrimental in medical AI, where models must generalize to novel modalities and diagnostic scenarios. Recent studies in general domains, including One-Shot RLVR (Wang et al. 2025), LIMO (Ye et al. 2025), and LIMR (Li, Zou, and Liu 2025), illustrate RL’s capacity for data-efficient learning and complex reasoning. In the medical domain, (Lai et al. 2025) and MedVLM-R1 (Pan et al. 2025) demonstrate that RL—specifically GRPO—substantially improves generalization when guided by high-quality preferences. However, these methods rely on the availability of curated reward datasets, a costly and unsustainable requirement in clinical contexts. Our approach addresses this shortfall by integrating a generative-reward pipeline that produces instruction samples and corresponding reward signals automatically.
3 Methodology
@@
We employ GRPO (Pan et al. 2025) to further refine $\pi_{\psi}$ under a composite reward that captures both reasoning quality and answer correctness. Specifically, the reward function is defined as:
@@
We use OmniMedVQA (Hu et al. 2024), a large-scale, multi-specialty medical VQA benchmark. It contains over 300k triplets covering radiology, dermatology, pathology, and ophthalmology, with diverse reasoning types (e.g., diagnosis, procedural, descriptive). We use the official train/val/test splits for all experiments.
@@
Our evaluation employs two distinct settings. First, for our main SOTA comparison (RQ1, RQ2), we follow the standard protocol of training models on the full, mixed-modality OmniMedVQA training set and evaluating their performance on the test set for each specific modality. Second, to specifically analyze knowledge transfer (RQ3), we adopt the stricter cross-domain protocol from Med-R1 (Lai et al. 2025), where models are trained on a single source domain (e.g., only MRI data) and evaluated on unseen target domains in a zero-shot manner.
@@
  • General-Purpose VLMs (Zero-shot): To assess the performance ceiling of large-scale vision-language models without any domain-specific tuning, we evaluate BLIP-2 (Li et al. 2023c), InstructBLIP (Dai et al. 2023), LLaVA (Li et al. 2024), MiniGPT-4 (Zhu et al. 2023), and LLaMA Adapter v2 (Gao et al. 2023), as well as strong recent models like Qwen2-VL-7B and Qwen2-VL-72B (Wang et al. 2024), and their upgraded versions Qwen2.5-VL-7B/72B (Bai et al. 2025).
  • Medical VLMs (Zero-shot): We include specialized models trained on biomedical data, such as LLaVA-Med (Li et al. 2023b), RadFM (Wu et al. 2023), Med-Flamingo (Moor et al. 2023), and MedVInT (Zhang et al. 2024), to benchmark domain-specific zero-shot capabilities.
@@
    • SFT Baselines: We fine-tune strong general-purpose models such as Qwen2.5-VL-3B and Qwen2.5-VL-7B using standard supervised instruction tuning. We also consider curated data generation pipelines like HuatuoGPT-Vision (Chen et al. 2024) and HealthGPT-Vision(Lin et al. 2025) when available.
    • RL Baselines: We include Med-R1 (Lai et al. 2025) variants (2B and 3B, with and without reasoning modules) as state-of-the-art reward-learning based VLMs, and compare directly against our own model variants: MedGR2 (SFT) and MedGR2 (SFT+GRPO).
@@
5 Conclusion
In this work, we introduced MedGR2, a novel framework that addresses the fundamental data scarcity and generalization challenges in medical VLM by creating a self-improving virtuous cycle of data generation and reward modeling. Our experiments show that MedGR2 achieves new state-of-the-art performance on the OmniMedVQA benchmark, outperforming strong SFT and RL baselines, and even substantially larger foundation models, demonstrating remarkable parameter efficiency. By transforming the core problem from one of data curation to one of intelligent, automated data generation, MedGR2 paves the way for developing more robust, generalizable, and scalable AI for high-stakes clinical reasoning.

### v2

@@ Abstract
The application of Vision-Language Models (VLMs) in medicine is critically hampered by the scarcity of high-quality, expert-annotated data. Supervised Fine-Tuning (SFT) on existing datasets often leads to poor generalization on unseen modalities and tasks, while Reinforcement Learning (RL), a promising alternative, is stymied by the lack of reliable reward signals in this data-scarce domain. To break this impasse, we introduce Generative Reward Learning for Medical Reasoning (MedGR2), a novel framework that creates a self-improving virtuous cycle. MedGR2 co-develops a data generator and a reward model, enabling the automated, continuous creation of high-quality, multimodal medical data that serves as a superior training source for the post-training. Our experiments demonstrate that SFT with MedGR2-produced data already surpasses baselines trained on large-scale, human-curated datasets. Crucially, when leveraging this data for RL via Group Relative Policy Optimization (GRPO), our model achieves state-of-the-art cross-modality and cross-task generalization, significantly outperforming specialized RL-based methods. Furthermore, our compact model, empowered by MedGR2, achieves performance competitive with foundation models possessing over 10 times more parameters. MedGR2 presents a new paradigm for data-efficient learning in high-stakes domains, transforming the problem from data scarcity to data generation and unlocking the full potential of RL for building truly generalizable medical AI.
Codehttps://github.com/HelmarZhi/MedGR-Square
Introduction
@@
Vision-language models (VLMs) hold great promise for transforming clinical reasoning, from multimodal diagnosis to radiology report generation. Their ability to jointly process and reason over medical images and texts offers an unprecedented opportunity to improve decision support, reduce clinician burden, and enable generalizable AI in high-stakes domains. Yet this vision is fundamentally bottlenecked by a key obstacle: the scarcity of high-quality, expert-annotated medical data (wang2025empowering; lu2025harnessing). Compared to general-domain tasks, acquiring medical data is constrained by privacy, annotation cost (guo2023information), and the domain-specific nature of expertise (10.1145/3737879), which results in narrow task coverage and limited generalization across unseen modalities and tasks.
@@
To address this challenge, the research community has explored two dominant paths, each with clear limitations. The first approach focuses on large-scale data curation and supervised fine-tuning (SFT). Some systems (chen2024towards) attempt to clean and reformat noisy biomedical corpora such as PubMed using large vision-language models (e.g., GPT-4o), producing instruction-style datasets for downstream training. While effective at increasing data scale, these methods remain fundamentally limited by the quality of their source material and the domain alignment of the models used for cleaning. More importantly, as recent studies highlight, SFT-trained models often memorize training distributions and struggle with generalization to new modalities and tasks. This constitutes a critical gap for real-world clinical deployment. (zhao2024robust).
@@
The second path is reinforcement learning (RL), which offers a more powerful paradigm for learning generalizable reasoning strategies. For instance, some works(lai2025med; pan2025medvlm) demonstrate that training with RL using small, expert-annotated preference datasets significantly improves cross-task and cross-modal performance over SFT baselines. However, RL in medicine faces a structural limitation: it relies on reward signals that are expensive to obtain, requiring human expert judgments for reward modeling or preference annotation. This creates a paradox: RL requires high-quality supervision to improve generalization, but such supervision is itself too costly to scale.
@@
Constructing high-quality, large-scale datasets remains a central bottleneck in developing medical VLMs. Manual annotation is prohibitively expensive and labor-intensive, prompting a surge of semi-automated solutions. HuatuoGPT-Vision (chen2024towards) reformats PubMed-derived image-text pairs into instruction-following VQA samples using GPT-4V. LLaVA-Med (jin2023llava) synthesizes multimodal QA pairs by prompting ChatGPT with biomedical figures. PMC-VQA (wang2023pmc) leverages large language models and rule-based heuristics to mine high-quality triplets from scientific publications. Flamingo-CXR (tanno2025collaboration) demonstrates the potential of radiologist-validated datasets, yet the extensive human oversight required severely limits scalability. While these approaches offer valuable contributions, they remain constrained by source data noise, domain coverage gaps, and alignment inconsistencies. In contrast, our framework abandons static corpora altogether, employing a self-improving generative pipeline to synthesize clinically meaningful, task-aligned multimodal instruction data from scratch.
@@
To mitigate data scarcity, generative approaches such as GANs and diffusion models have been widely adopted for synthesizing realistic medical images (Yi2019GenerativeMI; chowdhury2022diffusion; 10.1145/3737879). While effective for augmenting training sets in pixel-level tasks such as classification or segmentation, these techniques fall short in producing semantically coherent, instruction-following multimodal pairs required for medical reasoning. State-of-the-art VLMs demand data that captures clinical logic, procedural knowledge, and image-language alignment. Our framework transcends pixel-level synthesis by generating complete reasoning tasks paired with reward-guided quality filtering, enabling scalable construction of instruction-grade datasets for medical VLMs.
@@
Although SFT has facilitated early progress in VLMs, it often encourages rote pattern learning over robust generalization (chu2025sft). This limitation is especially detrimental in medical AI, where models must generalize to novel modalities and diagnostic scenarios. Recent studies in general domains, including One-Shot RLVR (wang2025reinforcement), LIMO (ye2025limo), and LIMR (li2025limr), illustrate RL’s capacity for data-efficient learning and complex reasoning. In the medical domain, (lai2025med) and MedVLM-R1 (pan2025medvlm) demonstrate that RL—specifically GRPO—substantially improves generalization when guided by high-quality preferences. However, these methods rely on the availability of curated reward datasets, a costly and unsustainable requirement in clinical contexts. Our approach addresses this shortfall by integrating a generative-reward pipeline that produces instruction samples and corresponding reward signals automatically.
Methodology
@@
We employ GRPO (pan2025medvlm) to further refine $\pi_{\psi}$ under a composite reward that captures both reasoning quality and answer correctness. Specifically, the reward function is defined as:
@@
We use OmniMedVQA (hu2024omnimedvqa), a large-scale, multi-specialty medical VQA benchmark. It contains over 300k triplets covering radiology, dermatology, pathology, and ophthalmology, with diverse reasoning types (e.g., diagnosis, procedural, descriptive). We use the official train/val/test splits for all experiments.
@@
Our evaluation employs two distinct settings. First, for our main SOTA comparison (RQ1, RQ2), we follow the standard protocol of training models on the full, mixed-modality OmniMedVQA training set and evaluating their performance on the test set for each specific modality. Second, to specifically analyze knowledge transfer (RQ3), we adopt the stricter cross-domain protocol from Med-R1 (lai2025med), where models are trained on a single source domain (e.g., only MRI data) and evaluated on unseen target domains in a zero-shot manner.
@@
  • General-Purpose VLMs (Zero-shot): To assess the performance ceiling of large-scale vision-language models without any domain-specific tuning, we evaluate BLIP-2 (li2023blip), InstructBLIP (dai2023instructblip), LLaVA (li2024llava), MiniGPT-4 (zhu2023minigpt), and LLaMA Adapter v2 (gao2023llama), as well as strong recent models like Qwen2-VL-7B and Qwen2-VL-72B (wang2024qwen2), and their upgraded versions Qwen2.5-VL-7B/72B (bai2025qwen2).
  • Medical VLMs (Zero-shot): We include specialized models trained on biomedical data, such as LLaVA-Med (li2023llava), RadFM (wu2023towards), Med-Flamingo (moor2023med), and MedVInT (zhang2024development), to benchmark domain-specific zero-shot capabilities.
@@
    • SFT Baselines: We fine-tune strong general-purpose models such as Qwen2.5-VL-3B and Qwen2.5-VL-7B using standard supervised instruction tuning. We also consider curated data generation pipelines like HuatuoGPT-Vision (chen2024towards) and HealthGPT-Vision(lin2025healthgpt) when available.
    • RL Baselines: We include Med-R1 (lai2025med) variants (2B and 3B, with and without reasoning modules) as state-of-the-art reward-learning based VLMs, and compare directly against our own model variants: MedGR2 (SFT) and MedGR2 (SFT+GRPO).
@@
Conclusion
In this work, we introduced MedGR2, a novel framework addressing data scarcity and generalization challenges in medical VLM via a self-improving virtuous cycle of data generation and reward modeling. Our experiments show MedGR2 achieves new state-of-the-art performance on the OmniMedVQA benchmark, outperforming strong SFT and RL baselines, and even substantially larger foundation models, demonstrating remarkable parameter efficiency. By transforming the core problem from data curation to intelligent, automated data generation, MedGR2 paves the way for more robust, generalizable, and scalable AI for high-stakes clinical reasoning.
Acknowledgments
This research was Supported by Guangdong S&T Program under Grant 2023B0303040002, and the Young Scientists Fund of the National Natural Science Foundation of China (C-Class, No. 32500997). The authors would like to express their sincere gratitude to Prof. Lirong Zheng and Prof. Yuxiang Huan for their constructive discussions and valuable feedback. The authors also thank all members of the lab for their continuous support.